OpenStudy (anonymous):

Bill scored 12 points in the first 8 minutes of the game. How many points would he score in the whole 40 minute game at this rate?

OpenStudy (anonymous):

320 or 27

hero (hero):

I calculated 60

OpenStudy (anonymous):

i was gonna say wouldnt he score more??? 60 seems about right

hero (hero):

12/8 = x/40 480 = 8x x = 60